# Reliable Projection Based Unsupervised Learning for Semi-Definite QCQP with Application of Beamforming Optimization

### v1

@@ I Introduction
With the development of 6G networks towards service-driven paradigms that prioritize user preferences, the optimization process in these networks faces an inevitable challenge of solving Constraint Optimization (CO) problems based on personalized user preferences and service performance constraints. While traditional iterative methods have demonstrated success in solving problems with precise convex constraints or providing approximate solutions for simple non-convex constraints, their iterative process often involves computationally intensive operations such as computing Hessian matrices or inverses. As a result, they fail to address real-time CO problems that require timely solutions, such as beamforming design issues with user rate preference constraints. This limitation undermines their applicability in time-sensitive applications. To overcome these challenges, researchers have turned to computationally efficient deep neural networks (NN), employing the Learning to Optimize (L2O) approach for solving CO problems. However, the utilization of L2O is not without flaws. The black-box nature of NNs, their data dependencies, deterministic parameters, and scaling law pose ongoing challenges in terms of efficiency, constraint handling, performance, and training costs that researchers cannot overlook
@@
While the universal approximation theorem suggests that NNs have the capability to model any mapping, thus theoretically implying the feasibility of designing an NN that can project input optimization problems into a feasible solution space, this assertion assumes the use of infinitely wide networks, which is impractical in real-world wireless network systems. When employing finite-width NNs, mapping errors are inevitable, leading to outputs that lie outside the feasible region. Consequently, existing methods that utilize L2O for solving CO problems often struggle to achieve satisfactory performance while maintaining low computational complexity. Even conventional approaches, such as constructing penalty functions associated with constraints to solve CO problems through unconstrained Lagrangian dual functions, fail to ensure computational efficiency. This is due to the increased complexity of the Lagrangian dual functions compared to the original CO problem’s objective functions, which in turn makes it more challenging to fit NNs. Moreover, these methods do not guarantee the feasibility of constraints as non-convex Lagrangian dual functions may lead NN to fall into local minima outside the feasible region. Existing feasible L2O methods, such as those based on sampling, require repetitive sampling within a fixed sampling space until a feasible point is found for the final output. However, this approach is computationally inefficient and limited to fixed constraint equations, unable to handle variable constraints. Such limitations contradict the diverse service requirements and personalized development goals of 6G networks, where user demands are dynamic and not always identical.
@@
Hence, it seems to increasingly necessary to enhance the model’s representational capacity to ensure the reliability of mappings from optimization problems to feasible solution spaces. As a result, NN architectures employed for wireless network beamforming optimization have evolved into more complex structures, transitioning from simple Multilayer Perceptrons (MLPs) to Convolutional NNs (CNNs), Long Short-Term Memory (LSTM) networks, and further to Graph NNs (GNNs) and Transformers. These advanced architectures possess unprecedented representational capabilities, enabling them to effectively tackle CO problems in radio frequency resource management. However, indiscriminate model size expansion offers minimal benefits for solving CO problems while introducing significant drawbacks. Larger models typically lead to longer inference delays, which conflicts with the trend of reducing the air interface transmission time interval (TTI) in the transition from 4G to 5G and toward future 6G networks. The reduction of TTI aims to enhance the agility and real-time capabilities of wireless radio frequency management. Consequently, the adoption of oversized models results in excessive inference delays, making their practical application in wireless network radio frequency resource management challenging. Furthermore, even when expanding the model parameters to the scale of hundreds of billions, there is no guarantee of producing constraint-compliant outputs. This is analogous to how even GPT-4 may occasionally generate sentences that do not adhere to grammatical and spelling rules.
@@ III-A Reliable Projection Function for Constrained MISO Beamforming
As Theorem 1 highlights, Problem 1 is NP-hard, suggesting that obtaining an optimal solution within polynomial time is unfeasible. Concurrently, the rapid advancements in NN technology provide a promising avenue for tackling NP-hard challenges. In the context of CO, employing NN-based strategies essentially involves training the network to approximate a function that maps the parameters of optimization and constraint equations directly to the optimal feasible solutions. The theoretical underpinning for this approach is supported by the universal approximation theorem, which asserts that a neural network with sufficient width can approximate any function to any desired degree of accuracy. However, practical constraints prevent the deployment of infinitely wide networks in real-world communication systems, leading inevitably to approximation errors as per the universal approximation theorem. Consequently, the output from a neural network may not always reside within the feasible region, nor can it guarantee inherently optimal solutions. Additionally, the opaque nature of neural networks complicates the analysis of conditions under which the network outputs points within the feasible region or the quality of the solutions it produces. To mitigate these issues, integrating a feature embedding module that compels the neural network to yield high-quality outputs within the feasible region becomes essential. Alternatively, employing a post-processing module to map potentially unreliable NN outputs back into the feasible region can ensure the performance quality of learning-to-optimize (L2O) schemes. Drawing inspiration from the successful application of softmax mapping in CO problems involving first-order linear sum constraints, a robust projection function has been devised, as depicted in Fig. LABEL:fig:system. This function is designed to facilitate a reliable CO solution leveraging neural network capabilities. The foundational design of this projection function is predicated on specific properties of Problems 1, as initially delineated.
@@ A. Individual Dataset Direct Execution
    Varying $M$: We set $N$ as 50, and $\gamma_{1}=\ldots=\gamma_{M}=\gamma=10$ dB. We consider different values of $M$, namely 40, 80, 120, 160, and 200. The simulation results are depicted in Fig. 3 and Fig. LABEL:fig:sqrt_user_time. The average loss for each model is depicted as bars, with accompanying black lines delineating the upper and lower performance bounds across different values of $M$. From the figures, it is evident that both Model 2 and Model 3 significantly outperform Model 1 and SDR with randomization in terms of average losses and training time. Model 1, which only projects infeasible solutions, exhibits weaker convergence compared to Models 2 and 3. Moreover, as the parameter $M$ increases, signifying a larger problem size, the Average Losses gradually increase. However, the impact on training time for projection-based methods is less pronounced than for SDR, suggesting a more scalable approach in terms of computational efficiency.
@@
    Varying $N$: We keep $M=80$ and $\gamma_{1}=\ldots=\gamma_{M}=\gamma=10$ dB, while varying $N$ with values of 10, 30, 50, 70, and 90 for experimentation. The results are depicted in Fig. 4 and Fig. LABEL:fig:sqrt_N_time. With increasing $N$, the problem size increases, but the difficulty of solving decreases. The average loss increases, the training time of the SDR increases, and the time of the projection-based methods decreases.
@@ B. Training on Large-Scale Datasets
    Varying $M$: Likewise, we retain the same configurations as previously described, with $N$ fixed at 50 and $\gamma_{1}=\ldots=\gamma_{M}=\gamma=10$ dB. We continue to examine various values of $M$, including 40, 80, 120, 160, and 200. From Fig. 6 and Fig. LABEL:fig:batch_user_time, it can be observed that the loss slightly increases with larger-scale datasets compared to single datasets, and the running time also increases. Meanwhile, the characteristic of Model 2 being difficult to converge is further magnified in larger-scale datasets.
@@
    Varying $N$: We maintain the same parameter settings as in the previous experiments, with $M=80$ and $\gamma_{1}=\ldots=\gamma_{M}=\gamma=10$ dB. The outcomes of these experiments are illustrated in Fig. 7 and Fig. LABEL:fig:batch_N_time.
@@
    Varying $\gamma_{i}$: We establish $N=50$ and $M=80$, while exploring a range of $\gamma$ values from 5dB to 25dB in increments of 5dB for our experimentation. The results are illustrated in Fig. 8 and Fig. LABEL:fig:batch_gamma_time.
@@ C. Model Deployment Strategies
In this section, we examine the time cost associated with deploying the model in practical applications. We fine-tune pre-trained models using large-scale datasets for inference in real-world scenarios. Employing function 5, we conduct inference on the new single dataset $h_{i}$ using the trained model, while varying the values of $M$, $N$, and $\gamma_{i}$. The runtime results are depicted in Figure LABEL:fig:all_three_images. From the figures, it is evident that projection-based learning optimization methods significantly outperform traditional SDR. They exhibit excellent timeliness and flexibility in practical applications.

### v2

@@ I Introduction
With the development of 6G networks towards service-driven paradigms that prioritize user preferences, the optimization process in these networks faces an inevitable challenge of solving constraint optimization (CO) problems based on personalized user preferences and service performance constraints [1]. While traditional iterative methods have demonstrated success in solving problems with precise convex constraints or providing approximate solutions for simple non-convex constraints, their iterative process often involves computationally intensive operations such as computing Hessian matrices or inverses [2, 3]. As a result, they fail to address real-time CO problems that require timely solutions, such as beamforming design issues with user rate preference constraints. This limitation undermines their applicability in time-sensitive applications. To overcome these challenges, researchers have turned to computationally efficient deep neural networks (NN), employing the learning to optimize (L2O) approach for solving CO problems. However, the utilization of L2O is not without flaws [4, 5]. The black-box nature of NNs, their data dependencies, deterministic parameters, and scaling law pose ongoing challenges in terms of efficiency, constraint handling, performance, and training costs that researchers cannot overlook
@@
While the universal approximation theorem suggests that NNs have the capability to model any mapping, thus theoretically implying the feasibility of designing an NN that can project input optimization problems into a feasible solution space, this assertion assumes the use of infinitely wide networks, which is impractical in real-world wireless network systems. When employing finite-width NNs, mapping errors are inevitable, leading to outputs that lie outside the feasible region [6]. Consequently, existing methods that utilize L2O for solving CO problems often struggle to achieve satisfactory performance while maintaining low computational complexity. Even conventional approaches, such as constructing penalty functions associated with constraints to solve CO problems through unconstrained Lagrangian dual functions, fail to ensure computational efficiency [7, 4, 5]. This is due to the increased complexity of the Lagrangian dual functions compared to the original CO problem’s objective functions, which in turn makes it more challenging to fit NNs. Moreover, these methods do not guarantee the feasibility of constraints as non-convex Lagrangian dual functions may lead NN to fall into local minima outside the feasible region. Existing feasible L2O methods, such as those based on sampling, require repetitive sampling within a fixed sampling space until a feasible point is found for the final output. However, this approach is computationally inefficient and limited to fixed constraint equations, unable to handle variable constraints. Such limitations contradict the diverse service requirements and personalized development goals of 6G networks, where user demands are dynamic and not always identical.
@@
Hence, it seems to increasingly necessary to enhance the model’s representational capacity to ensure the reliability of mappings from optimization problems to feasible solution spaces. As a result, NN architectures employed for wireless network beamforming optimization have evolved into more complex structures, transitioning from simple Multilayer Perceptrons (MLPs) [8] to Convolutional NNs (CNNs) [9], Long Short-Term Memory (LSTM) networks, and further to Graph NNs (GNNs) and Transformers [7, 10]. These advanced architectures possess unprecedented representational capabilities, enabling them to effectively tackle CO problems in radio frequency resource management. However, indiscriminate model size expansion offers minimal benefits for solving CO problems while introducing significant drawbacks. Larger models typically lead to longer inference delays, which conflicts with the trend of reducing the air interface transmission time interval (TTI) in the transition from 4G to 5G and toward future 6G networks. The reduction of TTI aims to enhance the agility and real-time capabilities of wireless radio frequency management [11]. Consequently, the adoption of oversized models results in excessive inference delays, making their practical application in wireless network radio frequency resource management challenging. Furthermore, even when expanding the model parameters to the scale of hundreds of billions, there is no guarantee of producing constraint-compliant outputs. This is analogous to how even GPT-4 may occasionally generate sentences that do not adhere to grammatical and spelling rules.
@@ III-A Reliable Projection Function for Constrained MISO Beamforming
As Theorem 1 highlights, Problem 1 is NP-hard, suggesting that obtaining an optimal solution within polynomial time is unfeasible. Concurrently, the rapid advancements in NN technology provide a promising avenue for tackling NP-hard challenges. In the context of CO, employing NN-based strategies essentially involves training the network to approximate a function that maps the parameters of optimization and constraint equations directly to the optimal feasible solutions. The theoretical underpinning for this approach is supported by the universal approximation theorem, which asserts that a neural network with sufficient width can approximate any function to any desired degree of accuracy. However, practical constraints prevent the deployment of infinitely wide networks in real-world communication systems, leading inevitably to approximation errors as per the universal approximation theorem. Consequently, the output from a neural network may not always reside within the feasible region, nor can it guarantee inherently optimal solutions. Additionally, the opaque nature of neural networks complicates the analysis of conditions under which the network outputs points within the feasible region or the quality of the solutions it produces. To mitigate these issues, integrating a feature embedding module that compels the neural network to yield high-quality outputs within the feasible region becomes essential. Alternatively, employing a post-processing module to map potentially unreliable NN outputs back into the feasible region can ensure the performance quality of learning-to-optimize (L2O) schemes. Drawing inspiration from the successful application of softmax mapping in CO problems involving first-order linear sum constraints, a robust projection function has been devised, as depicted in Fig. 1. This function is designed to facilitate a reliable CO solution leveraging neural network capabilities. The foundational design of this projection function is predicated on specific properties of Problems 1, as initially delineated.
@@ A. Individual Dataset Direct Execution
    Varying $M$: We set $N$ as 50, and $\gamma_{1}=\ldots=\gamma_{M}=\gamma=10$ dB. We consider different values of $M$, namely 40, 80, 120, 160, and 200. The simulation results are depicted in Fig. 3. The average loss for each model is depicted as bars, with accompanying black lines delineating the upper and lower performance bounds across different values of $M$. From the figures, it is evident that both Model 2 and Model 3 significantly outperform Model 1 and SDR with randomization in terms of average losses and training time. Model 1, which only projects infeasible solutions, exhibits weaker convergence compared to Models 2 and 3. Moreover, as the parameter $M$ increases, signifying a larger problem size, the Average Losses gradually increase. However, the impact on training time for projection-based methods is less pronounced than for SDR, suggesting a more scalable approach in terms of computational efficiency.
@@
    Varying $N$: We keep $M=80$ and $\gamma_{1}=\ldots=\gamma_{M}=\gamma=10$ dB, while varying $N$ with values of 10, 30, 50, 70, and 90 for experimentation. The results are depicted in Fig. 4. With increasing $N$, the problem size increases, but the difficulty of solving decreases. The average loss increases, the training time of the SDR increases, and the time of the projection-based methods decreases.
@@ B. Training on Large-Scale Datasets
    Varying $M$: Likewise, we retain the same configurations as previously described, with $N$ fixed at 50 and $\gamma_{1}=\ldots=\gamma_{M}=\gamma=10$ dB. We continue to examine various values of $M$, including 40, 80, 120, 160, and 200. From Fig. 6, it can be observed that the loss slightly increases with larger-scale datasets compared to single datasets, and the running time also increases. Meanwhile, the characteristic of Model 2 being difficult to converge is further magnified in larger-scale datasets.
@@
    Varying $N$: We maintain the same parameter settings as in the previous experiments, with $M=80$ and $\gamma_{1}=\ldots=\gamma_{M}=\gamma=10$ dB. The outcomes of these experiments are illustrated in Fig. 7.
@@
    Varying $\gamma_{i}$: We establish $N=50$ and $M=80$, while exploring a range of $\gamma$ values from 5dB to 25dB in increments of 5dB for our experimentation. The results are illustrated in Fig. 8.
@@ C. Model Deployment Strategies
In this section, we examine the time cost associated with deploying the model in practical applications. We fine-tune pre-trained models using large-scale datasets for inference in real-world scenarios. Employing function 5, we conduct inference on the new single dataset $h_{i}$ using the trained model, while varying the values of $M$, $N$, and $\gamma_{i}$. . From the figures, it is evident that projection-based learning optimization methods significantly outperform traditional SDR. They exhibit excellent timeliness and flexibility in practical applications.